OpenStudy (anonymous):

how do you graph 2x-5y=1 by using the x and y intercepts?

OpenStudy (texaschic101):

do you know how to find the x and y intercept ?

OpenStudy (anonymous):

no, I don't. this is really confusing.

OpenStudy (texaschic101):

to find the x intercept, sub on 0 for y. 2x - 5y = 1 2x - 5(0) = 1 2x = 1 x = 1/2 <-- x intercept (1/2,0) to find the y intercept, sub in 0 for x 2x - 5y = 1 2(0) - 5y = 1 -5y = 1 y = -1/5 <-- y intercept (0,-1/5) now plot your points and just draw a line and connect them.

OpenStudy (anonymous):

ok, thanks!

OpenStudy (texaschic101):

no problem :)